1/3

Дмитрий Николаевич Садовников, Дмитрий Николаевич Садовников, Дмитрий Николаевич Садовников,

Volga, Volga (Stenka Rasin)

Из-за острова на стрежень (Стенька Разин)

Sten’ka Razin

tradukita de N. N. 20  tradukita de Tr. Jacob Robbins

Bag om øen, på den brede underskønne Volgaflod Stenjka Rasins stolte både ud på nye togter stod.

1. Из-за острова на стрежень,

На простор речной волны, Выплывают расписные, Острогрудые челны.

1. From beyond the wooded island To the river wide and free Proudly sailed the arrow-breasted Ships of Cossack yeomanry.

På den første høvding Sten-

jka sidder selv i bådens stavn, han på kamp og strid ej

tænker, har prinsessen i sin favn.

2. На переднем Стенька Разин,

Обнявшись, сидит с

княжной, Свадьбу новую справляет, Сам веселый и хмельной.

2. On the first is Stenka Razin With a princess at his side, |: Drunken, holds a marriage revel With his beautiful young bride. :|

Men hans mænd er Donko-

sakker, de er vrede, for de ved, han har glemt dem for en

kvinde, for en nat i kærlighed.

3. А она, потупив очи, Ни жива и ни мертва, Молча слушает хмельные Атамановы слова.

3. But behind them rose a whisper, “He has left his sword to woo; |: One short night, and Stenka Ra-

zin Has become a woman too!” :|

Stenjka Rasin hører latter, blodet koger i hans bryst Over Volgas brede strømme højt nu lyder Stenjkas røst.

4. Позади их слышен ропот:

- Нас на бабу променял! Только ночь с ней

провозилс Сам наутро бабой стал .

4. Stenka Razin hears the jeering Of his discontented band, |: And the lovely Persian princess He has circled with his hand. :|

....................................... ....................................... ....................................... .......................................

5. Этот ропот и насмешки Слышит грозный атаман, И могучею рукою Обнял персиянки стан.

5. His black brows have come together

As the waves of anger rise,

|: And the blood comes rushing

swiftly To his piercing, jet-black eyes. :|

....................................... ....................................... ....................................... .......................................

6. Брови черные сошлися, Надвигается гроза. Буйной кровью налилис Атамановы глаза.

6. “I will give you all you ask for, Life and heart, and head and

hand,” |: Echo rolls the pealing thunder Of his voice across the land. :|

…  …  …

2/3

....................................... ....................................... ....................................... .......................................

7. “Ничего не пожалею, Буйну голову отдам!” — Раздается голос властный По окрестным берегам.

7. “Volga, Volga, mother Volga, Deep and wide beneath the sun, |: You have never seen a present From the Cossack of the Don. :|

....................................... ....................................... ....................................... .......................................

8. ”Волга, Волга, мать родная,

Волга, русская река, Не видала ты подарка От донского казака!

8. And that peace might rule as always

All my free-born men and brave, |: Volga, Volga, mother Volga, Volga, make this girl a grave.“ :|

”For at ingen splid skal være mellem frie mænd,- tag så Moder Volga, denne kvin-

de!“ og hans arm om hende lå.

9. Чтобы не было раздора Между вольными людьми, Волга, Волга, мать родная, На, красавицу возьми!”

9. With a sudden, mighty movement,

Razin lifts the beauty high,

|: And he casts her where the wa-

ters Of the Volga move and sigh. :|

Og han ta’r med stærke ar-

me da sin brud og kaster ned /: i de dybe, mørke bølger, sine nætters salighed. :/ 10. Мощным взмахом

поднимает Он красавицу княжну И за борт ее бросает В набежавшую волну.

10. Now a silence like the grave sinks

To all those who stand to see, |: And the battle-hardened Cos-

sacks Sink to weep on bended knee. :|

“Men hvad fanden, hvorfor

hænger I med næbbet? Bliv kun

ved! Dans og syng de røversange, syng for hendes sjælefred!”

11. “Что ж вы, братцы, приуныли?

Эй, ты, Филька, черт,

пляши! Грянем песню удалую На помин ее души!..”

11. “Dance, you fool, and men, make merry!

What has got into your eyes? |: Let us thunder out a chanty Of a place where beauty lies.” :|

Frem fra øen - ud på dybet stævner Stenjka Rasin ud /: med den stærke, stolte

både, sejler fra sin døde brud. :/ 12. Из-за острова на

стрежень, На простор речной волны, Выплывают расписные Острогрудые челны.

12. From beyond the wooded island To the river wide and free, |: Proudly sail the arrow-breasted Ships of Cossack yeomanry. :|

…  …  …

3/3

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Danan de N. N. 20.

Arg-478-1037 (2010-03-08 12:16:20)

Aperinte en 1947 en la kant-libro “DLEA-Kantaro” eldonita de Dana Laborista Esperanto-Asocio (DLEA), nro 100 kaj en la kantlibro “Store Esperanto-Sangbog”, SEFO 1939, n-ro 31. Vidu la retejon esperantonia. dk sub http: // www. esperantonia. dk/ volga. htm .

Verkinto de tiu ĉi Rusa poemo estas Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883).

Arg-478-927 (2010-03-07 18:59:55)

Pri la aŭtoro vidu ankaŭ la informojn sub http: // ru. wikipedia. org , la ruslingva parto de Wikipedia. Informoj pri la kanto, la melodio kaj diversaj versioj de la kanto estas troveblaj en la retejo http: // a-pesni. golosa. info/ popular20/ izzaostrova. htm .

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Anglan de Tr. Jacob Robbins en 1921.

Arg-478-1039 (2010-03-09 17:06:16)

Tiu ĉi angligo troviĝas en http: // ingeb. org/ songs/ stenkara. html .